sam (.sam.):

A problem Solve for x, 7.3^(x-1)+5-3^(x+1)=0

sam (.sam.):

\[\huge 7.3^{x-1}+5-3^{x+1}=0\]

OpenStudy (perl):

is that 7*3^(x-1) or (7.3)^(x-1)

sam (.sam.):

decimal

OpenStudy (perl):

hmmm, doesnt look easy

OpenStudy (lgbasallote):

do tell me..does this involve canceling 3^x-1 the solution i thought had that..

OpenStudy (anonymous):

\[3^x = \frac{15}{2}\] \[x = \log_3 \frac{15}2\]

sam (.sam.):

@Ishaan94 that's true if the decimal was "dot" *

OpenStudy (anonymous):

Oh, my bad... so it's \[(7.3)^{x-1} + 5-3^{x+1}=0\]Hmm Now I feel stupid :/

OpenStudy (perl):

ok I have an idea

OpenStudy (perl):